OpenStudy (phebe):

A line has slope -5/3 . Through which too points could this line pass? @Sparrowhawk69

OpenStudy (phebe):

want answer choices

OpenStudy (anonymous):

Yes please

OpenStudy (phebe):

ok

OpenStudy (phebe):

A:(12,13), (17,10) B: (16,15), (13,10) C: (0,7), (3,10) D: (11,13), (8,18)

OpenStudy (andrewkaiser333):

http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (andrewkaiser333):

you are doing point sslope form yes???

OpenStudy (phebe):

yes

OpenStudy (phebe):

but its a fraction

OpenStudy (andrewkaiser333):

uh huh

OpenStudy (phebe):

u made a mistake its a fraction k XD

OpenStudy (andrewkaiser333):

m=-5/3

OpenStudy (andrewkaiser333):

you plug it in

OpenStudy (phebe):

ok

OpenStudy (andrewkaiser333):

y – y1 = m(x – x1)

OpenStudy (andrewkaiser333):

that is formula

OpenStudy (phebe):

can u do it for me i don;'tt know how to use it

OpenStudy (phebe):

not the answer jsut plug it in

OpenStudy (andrewkaiser333):

y2-y1=m(x2-x1)

OpenStudy (phebe):

can u do it for me but not the answer k

OpenStudy (anonymous):

Pick any point on your graph. Say (0,0) ok?

OpenStudy (andrewkaiser333):

\[10-12=\frac{ -5 }{ 3}(17-13)\]

OpenStudy (phebe):

ok

OpenStudy (andrewkaiser333):

can you do the math for that?

OpenStudy (phebe):

is the answer A

OpenStudy (andrewkaiser333):

or you want me to go more

OpenStudy (phebe):

go aliol more

OpenStudy (anonymous):

Go down 5 units and over 3 to the point (3,-5) Go down 5 more unites and over 3 more from that point to (6, -10) These are two possible points a line with slope -5/3's can have.

OpenStudy (phebe):

k

OpenStudy (anonymous):

Refer to my answer for an easy method to determine your answer @Phebe.

OpenStudy (phebe):

ok

OpenStudy (andrewkaiser333):

ok \[-2=\frac{ -5 }{ 3 }(4)\]

OpenStudy (andrewkaiser333):

yeah just use his mine takes longer

OpenStudy (phebe):

so is the answer A

OpenStudy (andrewkaiser333):

mine does same thing

OpenStudy (phebe):

so u sayin its A

OpenStudy (phebe):

is it A @AndrewKaiser333

OpenStudy (andrewkaiser333):

i think so

OpenStudy (andrewkaiser333):

ask the other guy please just in case am trying to get my things answered as well

OpenStudy (anonymous):

I believe it's A

OpenStudy (phebe):

ok thahx

OpenStudy (phebe):

*thanx

OpenStudy (andrewkaiser333):

so i was right it is a

OpenStudy (phebe):

yes

OpenStudy (phebe):

no it was D

OpenStudy (anonymous):

Sorry

OpenStudy (phebe):

its ok

OpenStudy (andrewkaiser333):

sorry

OpenStudy (andrewkaiser333):

been a couple years since doing this subject part kinda rusty

OpenStudy (andrewkaiser333):

make sure to close it kk?